OpenStudy (jmansedg):

Solve: 8b + 20 ≥ 5b – 10. A. b ≤ 10/3 B. b ≤ 10 C. b ≥ -10 D. b ≥ -10/3

OpenStudy (cursedwizard):

8b - 5b = ? -10 - 20 = ?

OpenStudy (jmansedg):

-10-20= -30

OpenStudy (cursedwizard):

correct and 8b - 5b?

OpenStudy (jmansedg):

d?

OpenStudy (cursedwizard):

wrong

OpenStudy (jmansedg):

it is a right ?

OpenStudy (cursedwizard):

no

OpenStudy (jmansedg):

c?

OpenStudy (cursedwizard):

correct

OpenStudy (jmansedg):

Using the Pythagorean theorem, find the measure of the hypotenuse when a = 2.4in and b = 6.4in. A. 5.1in B. 5.8in C. 6.8in D. 7.9in

OpenStudy (cursedwizard):

\(2.4^2+6.4^2\)

OpenStudy (jmansedg):

a?

OpenStudy (cursedwizard):

wrong

OpenStudy (jmansedg):

sooooo

Mehek (mehek14):

Nice try @CursedWizard :)

OpenStudy (cursedwizard):

Thanks ^_^ I tried but they weren't cooperating :/